# Conformalized Polynomial Chaos Expansion for Uncertainty-aware Surrogate Modeling

### v1

@@ 2.3 Conformalized polynomial chaos expansion
where $h_{mm}\in\mathbb{R}$ is the $m$-th diagonal term of $\mathbf{H}$. The derivation of the analytical expression (14) is shown in appendix A.
For the predictive interval (12) based on the jackknife+, the LOO predictions $\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)\right\}_{m=1}^{M}$ for the test input $\mathbf{x}^{*}$ are additionally needed. Let us denote with $\mathbf{d}_{m}^{\top}\in\mathbb{R}^{1\times K}$ the $m$-th row of the design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, which contains the evaluation of the PCE basis upon the training input $\mathbf{x}^{(m)}$, such that $\mathbf{d}_{m}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{(m)}\right),\dots,\Psi_{K}\left(\mathbf{x}^{(m)}\right)\right)$. Accordingly, let us denote with $\mathbf{d}_{*}^{\top}\in\mathbb{R}^{1\times K}$ the vector of basis evaluations at the test input $\mathbf{x}^{*}$, such that $\mathbf{d}_{*}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{*}\right),\dots,\Psi_{K}\left(\mathbf{x}^{*}\right)\right)$. Then, the LOO predictions can be analytically computed as
@@
The derivation of the analytical expression (2.3) is shown in appendix B.
The analytical expressions (14) and (2.3) enable the efficient computation of jackknife and jackknife+ predictive intervals for PCE-based regression models, without the need for repeated model retraining. Note that these closed-form relations are valid for PCEs with a fixed polynomial basis. Should the basis be adaptively constructed, for example by means of sparsity-promoting algorithms [41, 42], explicit model refitting is required during the basis selection procedure. However, once the final basis is available, the analytical relations remain applicable.
@@ 2.4 Normalized non-conformity scores
In the previous section, the non-conformity scores were equal to the absolute LOO residuals, i.e., $\alpha_{m}=\left|r_{m}^{\text{LOO}}\right|$, $m=1,\dots,M$. Using the relative LOO error, defined as
@@
where $\mathrm{Var(Y)}$ denotes the variance of the output, normalized LOO-residual-based non-conformity scores can be defined as
@@ 3.  Numerical experiments
The conformalized PCE is applied to benchmark models that are commonly used for the validation of surrogate modeling methods. A univariate model is used in section 3.1, while multivariate models with six, seven, and ten input dimensions are used in sections 3.2, 3.3, and 3.4, respectively. The multivariate models are taken from the “Virtual Library of Simulation Experiments: Test Functions and Datasets” by Surjanovic and Bingham [50]. Both the jackknife and jackknife+ methods are employed, with either normalized or non-normalized non-conformity scores. Additionally, we investigate the capability of the conformalized PCE method to construct predictive intervals that provide the target coverage, depending on training data availability. In all cases, the experimental design used to train the PCE consists of samples drawn randomly from the distribution of the input parameters. After model training and predictive interval estimation, the conformalized PCE is compared against the true function based on a separate test dataset. The same test dataset is employed to estimate the empirical coverage. Each experiment is repeated for $100$ random seeds to estimate coverage and predictive interval statistics. For all test-cases, the target coverage is fixed at $95\%$ and the test dataset consists of $10^{4}$ random samples. A discussion based on the numerical results is available in section 3.5.
3.1 Meromorphic function
@@
with $a=1$ and $b=0.5$. We assume that the input $x$ takes values as uniformly distributed in $\left[-1,1\right]$. The conformalized PCE is constructed with maximum polynomial degrees $P\in\left\{2,3\right\}$. Since the input is one-dimensional, the polynomial basis consists of $P+1$ terms. The experimental design has size $M=C\left(P+1\right)^{2}$, where $C\in\left\{2,3,5,10\right\}$ is an oversampling coefficient. Note that the quadratic scaling of the experimental design in relation to the polynomial basis is necessary to ensure a well conditioned least squares problem (4) in the one-dimensional input case [39].
@@ 3.2 OTL circuit function
where the the input vector $\mathbf{x}\in\mathbb{R}^{6}$ consists of the five resistances $R_{b1}$, $R_{b2}$, $R_{f}$, $R_{c1}$, $R_{c2}$, and the current gain $\beta$. The response corresponds to the midpoint voltage. The input parameters are assumed to be uniformly distributed within the value ranges given in Table 1.
@@
Total-degree conformalized PCEs with maximum polynomial degrees $P\in\left\{1,2,3\right\}$ and oversampling coefficients $C\in\left\{2,3,5,10\right\}$ are employed. Following Migliorati et al. [39], the experimental design now scales linearly with the size of the PCE basis, such that $M=CK=C\#\Lambda$. The different conformalized PCE configurations are compared against the true OTL circuit function for a single random seed in Figure 4. This plot does not compare normalized versus non-normalized conformity scores, as the differences between the corresponding predictive intervals are not easily discernible. Figures 5 and 6 shows box plots regarding the distributions of the conformalized PCEs’ empirical coverages and predictive intervals over the 100 random seeds, also including the two different non-conformity scores.
@@
3.3 Piston simulation function
@@
$\text{where }V=\frac{S}{2k}\left(\sqrt{A^{2}+4k\frac{P_{0}V_{0}}{T_{0}}T_{a}}-A\right)\text{ and }A=P_{0}\,S+19.62\,M-\frac{k\,V_{0}}{S},$
where the input vector $\mathbf{x}\in\mathbb{R}^{7}$ consists of the parameters listed in Table 2 along with their value ranges. All parameters are assumed to be uniformly distributed within their ranges.
@@
Total-degree conformalized PCEs with maximum polynomial degrees $P\in\left\{2,3,4\right\}$ and oversampling coefficients $C\in\left\{2,3,5,10\right\}$ are employed. Similar to section 3.2, the experimental design scales linearly with the size of the PCE basis, i.e., $M=CK=C\#\Lambda$. Figure 7 compares the various configurations of the conformalized PCE against the true piston simulation function for a single random seed. Figures 8 and 9 show coverage and predictive interval statistics over the 100 random seeds.
@@
3.4 Wing weight function
@@
3.5 Discussion on the numerical results
The numerical results presented in sections 3.1-3.4 show a number of patterns and similarities. First, the conformalized PCE is in general capable of providing empirical coverages close to the target level, more commonly above it. For less accurate PCEs, e.g., ones corresponding to low polynomial degrees and/or small experimental designs, the coverage over the 100 random seeds varies significantly. As the accuracy of the PCE increases with larger experimental designs and higher polynomial degrees, the variability in coverage is reduced and the coverage moves closer to the target level.
The coverage provided by the conformalized PCE is naturally connected to the corresponding predictive intervals. For the less accurate PCEs, the required coverage is accomplished by larger predictive intervals, which also show a large variability over the 100 random seeds. As expected, the predictive intervals become smaller as the accuracy of the PCE increases. Importantly, even for low-degree PCEs, increasing the size of the experimental design leads not only to a reduced variability in the predictive interval, but also to a reduced interval width.
@@ 3.5 Discussion on the numerical results
Last, comparing normalized and non-normalized non-conformity scores, no concrete conclusions can be made based on the available numerical results. There exist differences between the two, which are more common for PCEs trained with small experimental designs. However, these differences are generally small and not consistent. Overall, it can be said that both non-conformity scores perform similarly.

### v2

@@ 2.3 Conformalized polynomial chaos regression
where $h_{mm}\in\mathbb{R}$ is the $m$-th diagonal term of $\mathbf{H}$. The derivation of the analytical expression (14) is given in appendix A.
For the predictive interval (12) based on the jackknife+, the LOO predictions $\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)\right\}_{m=1}^{M}$ for the test input $\mathbf{x}^{*}$ are additionally needed. Let us denote with $\mathbf{d}_{m}^{\top}\in\mathbb{R}^{1\times K}$ the $m$-th row of the design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, which contains the evaluation of the PCE basis on the training input $\mathbf{x}^{(m)}$, such that $\mathbf{d}_{m}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{(m)}\right),\dots,\Psi_{K}\left(\mathbf{x}^{(m)}\right)\right)$. Accordingly, let us denote with $\mathbf{d}_{*}^{\top}\in\mathbb{R}^{1\times K}$ the vector of basis evaluations at the test input $\mathbf{x}^{*}$, such that $\mathbf{d}_{*}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{*}\right),\dots,\Psi_{K}\left(\mathbf{x}^{*}\right)\right)$. Then, the LOO predictions can be analytically computed as
@@
The derivation of the analytical expression (2.3) is given in appendix B.
The analytical expressions (14) and (2.3) enable the efficient computation of jackknife and jackknife+ predictive intervals for PCE-based regression models, without the need for repeated model retraining. Note that these closed-form relations are valid for PCEs with fixed polynomial basis. Should the basis be adaptively constructed, for example by means of sparsity-promoting algorithms [luethen2021sparse, luethen2022automatic], explicit model refitting is required during the basis selection procedure. However, once the final basis is available, the analytical expressions remain applicable.
@@ 2.4 Normalized non-conformity scores
In the previous section, the non-conformity scores were equal to the absolute LOO residuals, i.e., $\alpha_{m}=\left|r_{m}^{\text{LOO}}\right|$, $m=1,\dots,M$. Based on the relative LOO error, defined as
@@
where $\mathrm{Var(Y)}$ denotes the variance of the output, normalized non-conformity scores can be defined as
@@ 3.  Numerical experiments
The conformalized PCE method is first applied to four benchmark models that are commonly used for the validation of surrogate modeling methods (section 3.1). Then, the method is applied to two electrical engineering use-cases concerning the design of, first, a power module heat sink (section 3.2) and, second, a Stern-Gerlach electromagnet (section 3.3). Different conformalized PCE configurations are examined, depending on jackknife method, maximum polynomial degree, non-conformity score type and, most importantly, training data availability. The conformalized PCE models are assessed with respect to their ability to construct predictive intervals that provide the required coverage. Separate test datasets are used for that purpose. Each experiment is repeated for $100$ random seeds to estimate coverage and predictive interval statistics.
3.1 Validation on benchmark models
A univariate model is used in section 3.1.1, while multivariate models with six, seven, and ten input dimensions are used in sections 3.1.2, 3.1.3, and 3.1.4, respectively. The multivariate models are taken from the “Virtual Library of Simulation Experiments: Test Functions and Datasets” by Surjanovic and Bingham [surjanovic2013virtual]. For all benchmark models, the experimental design used to train the PCE consists of samples drawn randomly from the distribution of the input parameters. The target coverage is fixed at $95\%$ and the test datasets consist of $10^{4}$ random samples. A discussion based on the numerical results obtained from the benchmark models is available in section 3.1.5.
3.1.1 Meromorphic function
@@
with $a=1$ and $b=0.5$. We assume that the input $x$ takes values as uniformly distributed in $\left[-1,1\right]$. The conformalized PCE is constructed with maximum polynomial degrees $P\in\left\{2,3\right\}$. Since the input is one-dimensional, the polynomial basis consists of $P+1$ terms. The experimental design has size $M=C\left(P+1\right)^{2}$, where $C\in\left\{2,3,5,10\right\}$ is an oversampling coefficient. Note that the quadratic scaling of the experimental design in relation to the polynomial basis is necessary to ensure a well conditioned least squares problem (4) in the one-dimensional input case [migliorati2014analysis].
@@ 3.1.2 OTL circuit function
where the input vector $\mathbf{x}\in\mathbb{R}^{6}$ consists of the resistances $R_{b1}$, $R_{b2}$, $R_{f}$, $R_{c1}$, $R_{c2}$, and the current gain $\beta$. The response corresponds to the midpoint voltage. The input parameters are assumed to be uniformly distributed within the value ranges given in Table 1.
@@
Total-degree conformalized PCEs with maximum polynomial degrees $P\in\left\{1,2,3\right\}$ and oversampling coefficients $C\in\left\{2,3,5,10\right\}$ are employed. Following Migliorati et al. [migliorati2014analysis], the experimental design now scales linearly with the size of the PCE basis, such that $M=CK=C\#\Lambda$. The different conformalized PCE configurations are compared against the true OTL circuit function for a single random seed in Figure 4. This plot does not compare normalized versus non-normalized conformity scores, as the differences between the corresponding predictive intervals are not easily discernible. Figures 5 and 6 shows box plots regarding the distributions of the conformalized PCEs’ empirical coverages and predictive intervals over the 100 random seeds, also including the two different non-conformity scores.
@@
3.1.3 Piston simulation function
@@
$\text{where }V=\frac{S}{2k}\left(\sqrt{A^{2}+4k\frac{P_{0}V_{0}}{T_{0}}T_{a}}-A\right)\text{ and }A=P_{0}\,S+19.62\,M-\frac{k\,V_{0}}{S}.$
The input vector $\mathbf{x}\in\mathbb{R}^{7}$ consists of the parameters listed in Table 2 along with their value ranges. All parameters are assumed to be uniformly distributed within their ranges.
@@
Total-degree conformalized PCEs with maximum polynomial degrees $P\in\left\{2,3,4\right\}$ and oversampling coefficients $C\in\left\{2,3,5,10\right\}$ are employed. Similar to section 3.1.2, the experimental design scales linearly with the size of the PCE basis, i.e., $M=CK=C\#\Lambda$. Figure 7 compares the various configurations of the conformalized PCE against the true piston simulation function for a single random seed. Figures 8 and 9 show coverage and predictive interval statistics over the 100 random seeds.
@@
3.1.4 Wing weight function
@@
3.1.5 Discussion on benchmarks’ results
The numerical results presented in sections 3.1.1-3.1.4 show a number of patterns and similarities. First, the conformalized PCE is in general capable of providing empirical coverages close to the target level, more commonly above it. For less accurate PCEs, e.g., ones corresponding to low polynomial degrees and/or small experimental designs, the coverage over the 100 random seeds varies significantly. As the accuracy of the PCE increases with larger experimental designs and higher polynomial degrees, the variability in coverage is reduced and the coverage moves closer to the target level. The size of the experimental design is particularly important, as a sufficiently training dataset enables even low-accuracy conformalized PCEs to reach the target coverages.
The coverage provided by the conformalized PCE is naturally connected to the corresponding predictive intervals. For the less accurate PCEs, the required coverage is accomplished by larger predictive intervals, which also show a large variability over the 100 random seeds. As expected, the predictive intervals become smaller as the accuracy of the PCE increases. Importantly, even for low-accuracy PCEs, increasing the size of the experimental design leads not only to a reduced variability in the predictive interval, but also to a reduced interval width.
@@ 3.1.5 Discussion on benchmarks’ results
Last, comparing normalized and non-normalized non-conformity scores, no concrete conclusions can be made based on the available numerical results. There exist differences between the two, which are more common for PCEs trained with small experimental designs. However, these differences are generally small and not consistent. Overall, it can be said that the two non-conformity scores perform similarly.
3.2 Application to power module heat sink design
[…]
(a) Power electronic module.
(b) Heat sink simulation model.
Figure 13: Illustration of the power module and of the CFD heat sink simulation model.
Table 4: Input parameters of the power module heat-sink model.
Parameter Description Unit Range
$l$ heat sink length $\mathrm{mm}$ $\left[50,200\right]$
$g_{\text{f}}$ fin gap $\mathrm{mm}$ $\left[3,8\right]$
$w_{\text{f}}$ fin width $\mathrm{mm}$ $\left[1.4,4\right]$
$h_{\text{f}}$ fin height $\mathrm{mm}$ $\left[16,45\right]$
$h_{\text{b}}$ base height $\mathrm{mm}$ $\left[4,15\right]$
$N_{\text{f}}$ number of fins $\left[5,25\right]$
$v$ air flow velocity $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ $\left[1,5\right]$
[…]
$P$ power loss $\mathrm{W}$ $\left[115,140\right]$
To enable computationally demanding design exploration studies such as heat sink design optimization [loukrezis2022power], it is necessary to replace the CFD model with an inexpensive surrogate model. Here we are concerned with the surrogate modeling task only, with additional interest in the predictive uncertainty of the surrogate. Two significance levels are considered, namely, $s=0.05$ and $s=0.01$, respectively corresponding to target coverage levels of 1s=0.951-s=0.95 and 1s=0.991-s=0.99. The conformalized PCE has maximum polynomial degree $P\in\left\{2,3\right\}$. A dataset consisting of $935$ design configurations along with the corresponding thermal resistance values is readily available. Experimental designs of increasing size $M\in\left\{250,300,\dots,500\right\}$ are employed, while $M^{\prime}=435$ data points are kept as test data. The initial dataset is reshuffled $100$ times to estimate coverage and predictive interval statistics over the different dataset partitions.
(a) $P=2$, $M=300$, $s=0.05$.
(b) $P=3$, $M=300$, $s=0.05$.
(c) $P=2$, $M=300$, $s=0.01$.
(d) $P=3$, $M=300$, $s=0.01$.
(e) $P=2$, $M=400$, $s=0.05$.
(f) $P=3$, $M=400$, $s=0.05$.
(g) $P=2$, $M=400$, $s=0.01$.
(h) $P=3$, $M=400$, $s=0.01$.
(i) $P=2$, $M=500$, $s=0.05$.
(j) $P=3$, $M=500$, $s=0.05$.
(k) $P=2$, $M=500$, $s=0.01$.
(l) $P=3$, $M=500$, $s=0.01$.
Figure 14: Parity plots comparing ground truth $R_{\text{th,s}}$ values given by the heat-sink simulation model against conformalized PCE predictions for different combinations of polynomial degree $P$, training dataset size $M$ and significance level $s$. The results correspond to a single partition of the available design dataset. The results with and without non-conformity score normalization are very similar, therefore, only one set of results in shown.
(a) $P=2$, $\alpha_{m}$, $s=0.05$.
(b) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(c) $P=3$, $\alpha_{m}$, $s=0.05$.
(d) $P=3$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(e) $P=2$, $\alpha_{m}$, $s=0.01$.
(f) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
(g) $P=3$, $\alpha_{m}$, $s=0.01$.
(h) $P=3$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
Figure 15: Box plots of the empirical coverage provided by conformalized PCE surrogates of the heat-sink model, for different combinations of polynomial degree $P$, training dataset size $M$, significance level $s$, and non-conformity score type.
(a) $P=2$, $\alpha_{m}$, $s=0.05$.
(b) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(c) $P=3$, $\alpha_{m}$, $s=0.05$.
(d) $P=3$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(e) $P=2$, $\alpha_{m}$, $s=0.01$.
(f) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
(g) $P=3$, $\alpha_{m}$, $s=0.01$.
(h) $P=3$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
Figure 16: Box plots of the predictive interval widths provided by conformalized PCE surrogates of the heat-sink model, for different combinations of training dataset size $M$, significance level $s$, and non-conformity score type.
Figure 14 compares $R_{\text{th,s}}$ predictions provided by the conformalized PCE against the true CFD model outputs for training dataset sizes $M\in\left\{300,400,500\right\}$. Only the original dataset is used for these plots, without data reshuffling. Figures 15 and 16 show coverage and predictive interval statistics over the $100$ different dataset partitions and for all considered training dataset sizes. The empirical coverages provided by the conformalized PCEs are typically on par or slightly above the target levels. Jackknife CP is again more conservative compared to jackknife+, especially for the more demanding significance level $s=0.01$. Nonetheless, both approaches perform very similarly as the training dataset size increases. This is also reflected on the predictive interval widths, where, as would be expected, the significance level $s=0.01$ results in wider interval widths. The use of normalized non-conformity scores results in differences regarding coverage and predictive interval width. However, there is no consistent benefit or drawback in using one instead of the other.
3.3 Application to electromagnet design
[…]
The target model output is the average magnetic field gradient in the beam area, denoted as $\overline{\tau}_{\text{b}}$ and measured in $\mathrm{T}\text{\,}{\mathrm{m}}^{-1}$. For computational efficiency, $\overline{\tau}_{\text{b}}$ is computed with a 2D isogeometric analysis (IGA) model of the pole region, while the coils and yoke of the original 3D model are replaced by equivalent circuit models [pels2015optimization]. Then, $\overline{\tau}_{\text{b}}$ is estimated as
$\overline{\tau}_{\text{b}}=\frac{1}{\Omega_{\text{b}}}\int_{\Omega_{\text{b}}}\tau(x,y)\mathrm{d}\Omega,$ (24)
where $\Omega$ denotes the full 2D computational domain, $\Omega_{\text{b}}$ corresponds to the beam area, and $\tau(x,y)=\frac{\partial\left|\vec{B}\right|}{\partial x}$ is the magnetic field gradient in the $x$-direction, where $\vec{B}$ denotes the magnetic flux density. The latter is obtained by the IGA model’s field solution.
(a) Stern-Gerlach magnet model (half geometry).
(b) Pole region geometry.
Figure 17: Illustration of the Stern-Gerlach magnet model.
While much more efficient compared to the original 3D model, the circuit-coupled IGA model is still too computationally expensive for design optimization [masschaele2011design, pels2015optimization] or UQ [loukrezis2019assessing] studies, hence the need for a surrogate model. Here, the surrogate model is used to approximate $\overline{\tau}_{\text{b}}$ under random shape deformations in the pole region. The geometry of the pole region is described using non-uniform B-splines (NURBS) $\mathbf{R}\left(\xi\right)$, $\mathbf{\xi}\in\left[0,1\right]$, defined by control points $\mathbf{P}_{i}$, $i=1,\dots,N_{\text{cp}}$, given as
$\mathbf{R}\left(\xi\right)=\sum_{i=1}^{N_{\text{cp}}}\mathbf{P}_{i}\frac{w_{i}B_{i}^{p}\left(\xi\right)}{\sum_{j=1}^{N_{\text{cp}}}w_{j}B_{j}^{p}\left(\xi\right)},$ (25)
where $B_{i}^{p}\left(\xi\right)$ are B-spline basis functions of order $p$ and $w_{i}$ are their corresponding weights. Random shape deformations are introduced by modeling the coordinates of five control points and four of the corresponding weights as uniform random variables. All input parameters and their ranges are listed in Table 5.
Table 5: Input parameters of the Stern-Gerlach magnet model.
Parameter Description Unit Range
$_{1} xx$-coordinate, control point 1 $\mathrm{mm}$ $\left[-3.38,-1.38\right]$
$_{1} yy$-coordinate, control point 1 $\mathrm{mm}$ $\left[5.96,7.96\right]$
$_{2} xx$-coordinate, control point 2 $\mathrm{mm}$ $\left[-3.38,-1.38\right]$
$_{2} yy$-coordinate, control point 2 $\mathrm{mm}$ $\left[3.96,5.96\right]$
$_{3} xx$-coordinate, control point 3 $\mathrm{mm}$ $\left[16,18\right]$
$_{3} yy$-coordinate, control point 3 $\mathrm{mm}$ $\left[19,21\right]$
$_{4} xx$-coordinate, control point 4 $\mathrm{mm}$ $\left[-18,-16\right]$
$_{4} yy$-coordinate, control point 4 $\mathrm{mm}$ $\left[19,21\right]$
$_{5} xx$-coordinate, control point 5 $\mathrm{mm}$ $\left[-7,-5\right]$
$_{5} yy$-coordinate, control point 5 $\mathrm{mm}$ $\left[3,2\right]$
$w_{1}$ weight, control point 1 $\left[0,1\right]$
$w_{2}$ weight, control point 2 $\left[0,1\right]$
$w_{3}$ weight, control point 3 $\left[0,1\right]$
$w_{4}$ weight, control point 4 $\left[0,1\right]$
Two significance levels $s=0.05$ and $s=0.01$ are considered regarding the predictive uncertainty of the surrogate model. The conformalized PCE is applied with maximum polynomial degrees $P\in\left\{1,2\right\}$. A dataset with $1000$ different shape deformations of the pole region is available, along with the corresponding $\overline{\tau}_{\text{b}}$ vaues. Training datasets of increasing size $M\in\left\{150,200,\dots,500\right\}$ are employed. The remaining $M^{\prime}=500$ data points are used for validation. The dataset is shuffled $100$ times to estimate coverage and predictive interval statistics.
(a) $P=1$, $M=150$, $s=0.05$.
(b) $P=2$, $M=150$, $s=0.05$.
(c) $P=1$, $M=150$, $s=0.01$.
(d) $P=2$, $M=150$, $s=0.01$.
(e) $P=1$, $M=300$, $s=0.05$.
(f) $P=2$, $M=300$, $s=0.05$.
(g) $P=1$, $M=300$, $s=0.01$.
(h) $P=2$, $M=300$, $s=0.01$.
(i) $P=1$, $M=450$, $s=0.05$.
(j) $P=2$, $M=450$, $s=0.05$.
(k) $P=1$, $M=450$, $s=0.01$.
(l) $P=2$, $M=450$, $s=0.01$.
Figure 18: Parity plots comparing ground truth $\overline{\tau}_{\text{b}}$ values given by the Stern-Gerlach magnet model against conformalized PCE predictions for different combinations of polynomial degree $P$, training dataset size $M$ and significance level $s$. The results correspond to a single partition of the available design dataset. The results with and without non-conformity score normalization are very similar, therefore, only one set of results in shown.
(a) $P=1$, $\alpha_{m}$, $s=0.05$.
(b) $P=1$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(c) $P=2$, $\alpha_{m}$, $s=0.05$.
(d) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(e) $P=1$, $\alpha_{m}$, $s=0.01$.
(f) $P=1$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
(g) $P=2$, $\alpha_{m}$, $s=0.01$.
(h) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
Figure 19: Box plots of the empirical coverage provided by conformalized PCE surrogates of the Stern-Gerlach magnet model, for different combinations of polynomial degree $P$, training dataset size $M$, significance level $s$, and non-conformity score type.
(a) $P=1$, $\alpha_{m}$, $s=0.05$.
(b) $P=1$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(c) $P=2$, $\alpha_{m}$, $s=0.05$.
(d) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.05$.
(e) $P=1$, $\alpha_{m}$, $s=0.01$.
(f) $P=1$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
(g) $P=2$, $\alpha_{m}$, $s=0.01$.
(h) $P=2$, $\alpha_{m}^{\text{norm}}$, $s=0.01$.
Figure 20: Box plots of the predictive interval widths provided by conformalized PCE surrogates of the Stern-Gerlach magnet model, for different combinations of training dataset size $M$, significance level $s$, and non-conformity score type.
Figure 18 compares ground truth $\overline{\tau}_{\text{b}}$ values provided by the IGA model against conformalized PCE predictions for the original dataset without reshuffling. Training dataset sizes $M\in\left\{150,300,450\right\}$ are considered in this figure. Figures 19 and 20 show boxplots that summarize the statistics concerning empirical coverage and predictive interval width over the $100$ dataset reshuffles and for all considered training dataset sizes. The conformalized PCE provides the required coverage, even for the inaccurate $P=1$ models and the more demanding significance level $s=0.01$. As would be expected, the predictive intervals become narrower for $P=2$, while still preserving the required coverage. Similar to the previous test-cases, the empirical coverages and the predictive interval widths present high variations for smaller experimental designs that progressively stabilize as more training data become available. The jackknife method is once again the more conservative option, while non-conformity score normalization does not seem to have a significant impact.